What is the mole fraction of ethanol in a solution made by dissolving 29.2 g of ethanol, C2H5OH, in 53.6 g of water?
Answer #1

moles of ethanol = 29.2 / 46 = 0.634

moles of water = 53.6 / 18 = 2.978

molefraction of ethanol =0.634/(0.634+2.978) = 0.175
Answer #2

Etanol : 29.2/46 =0.63

water : 53.6/18 =2.97

Mole fraction of ethanol = 0.63/(0.63+2.97)

= 0.63/3.6

= 0.175